I need help with this question
If 500 is deposited in an account paying 8.5% annual interest, compounded continuously, how long will it take for the account to increase to $800? calculate your answere to three decimal places before rounding to two.

Answer by stanbon(48569)

If 500 is deposited in an account paying 8.5% annual interest, compounded continuously, how long will it take for the account to increase to $800? calculate your answere to three decimal places before rounding to two.
--------
A(t) = A(o)e^(rt)
800 = 500*e^(0.085t)
e^(0.085t) = 8/5
Take the natural log of both sides to get:
0.085t = 0.47
t = 0.47/0.85
t = 5.53 years. (rounded to two decimal places)
==============
Cheers,
Stan H.